OpenStudy (anonymous):

Homework 12 and LAB12 answers......! H12P1: CURRENT SOURCE 1) 1000 2) 2200 3) 0.004926 H12P2 LINEAR REGULATOR 1) vin*(R1+R2)/R2 2) vz*(1+(R1/R2)) 3) 700 4) 500 5) 500 6) 0.15 H12P3: OPAMPS AND FILTER DESIGN 1) 23417 2) 21489 3) 20000 4) 6.8e-7 5) 3.9e-10 6) 23417 7) 21489 8) 6.8e-7 9) 3.9e-10 10) -0.917

OpenStudy (anonymous):

LAB 12 R1=10 R2=2490 C1=0.0000001 C2=0.0000001

OpenStudy (anonymous):

HWP3 filter: 10Hz and 19kHz. signals around 1kHz, to be amplified by about a factor of 2. 0.68μF and a 390pF capacitor. resistors between 100Ω and 100kΩ. RS=20kΩ. pi=3.14 part1) First find C1=0.68*10^-6 = 6.8e-7 (answer of part 4) Then R1*C1=1/(2*pi*f) R1*C1=1/(2*pi*10) so R1= 23417 part 2)First find C2=390*10^-12 = 3.9e-10 (answer of part 5) R2*C2=1/(2*pi*19*10^3) R2= 21489 part 3) Rs = Rf = 20kΩ = 20000 Ω part 6 : R3= R1 part 7 : R4= R2 part 8: c3=c1 part 9 : c4= c2 part 9 :Midband gain = ­(R4/R3)

OpenStudy (anonymous):

lap 12 not correct for me, please help

OpenStudy (anonymous):

any one did lab 12 please help how to calculate R1 C1 R2 C2